OpenStudy (anonymous):

Solve. Choose the letter for the best answer. (Must show work!) 1. 2/5a = 10 A. a = 5 B. a = 20 C. a = 25

hartnn (hartnn):

multiply 5 on both sides, what do u get ?

OpenStudy (anonymous):

8

hartnn (hartnn):

how ?

OpenStudy (anonymous):

by nutiplying both sides by five

hartnn (hartnn):

what do u get on left side ? 5* (2/5) x = ... ?

hartnn (hartnn):

sorry, its 'a' 5*(2/5) a = ... ?

OpenStudy (anonymous):

125

hartnn (hartnn):

no... the 5's cancel out and you just get 2a on left

hartnn (hartnn):

2a = 10*5 = 50 a = 50/2

OpenStudy (anonymous):

i need the work shown

OpenStudy (anonymous):

so i can understand it

hartnn (hartnn):

did you not understand my steps ? what doubt u have ?

OpenStudy (anonymous):

i did

OpenStudy (math&ing001):

Must so frustrating @hartnn

OpenStudy (arbershabani97):

2/5a=10 /x5 2a=50 //2 a=25

hartnn (hartnn):

not actually, i've had worse

OpenStudy (math&ing001):

This is why sometimes I just give the answer before I get pissed off.

hartnn (hartnn):

in any case, giving out answer directly is not allowed at all

OpenStudy (anonymous):

It's hard to say from the original syntax, but it looks like you're saying (2/5)a=10. If so, you need to isolate a by turning the coefficient in front of a into 1. Do that by multiplying it by the reciprocal (5/2). In doing so, you must also do the same to the other side of the equation. Therefore: (5/2)(2/5)a = 10(5/2) Leaving you with a=10(5/2) This is the same as 50/2, or, 25.